# Persistent Homology via Finite Topological Spaces

### v1

@@ Abstract
We propose a functorial framework for persistent homology based on finite topological spaces and their associated posets. Starting from a finite metric space, we associate a filtration of finite topologies whose structure maps are continuous identity maps. By passing functorially to posets and to simplicial complexes via crosscut constructions, we obtain persistence modules without requiring inclusion relations between the resulting complexes. We show that standard poset-level simplifications preserve persistent invariants and prove stability of the resulting persistence diagrams under perturbations of the input metric in a density-based instantiation.
@@
Keywords: Persistent homology; finite topological spaces; posets; crosscut complexes; stability.
@@ 1 Introduction
In this paper, we develop a functorial framework for persistent homology that interpolates between metric data and simplicial complexes by way of finite topological spaces and posets. Starting from a finite metric space, we first associate to it a family of finite topologies indexed by a scale parameter. These topologies encode relational or density-based information and form a filtration via continuous identity maps. By passing functorially to associated posets and then to simplicial complexes via crosscut constructions, we obtain persistence modules without requiring inclusion relations between the resulting complexes.
This perspective separates the role of the metric as a generator of topological structure from the subsequent homological analysis. As a result, it allows for topological and combinatorial simplifications at the poset level, while preserving persistent invariants. Moreover, stability of the resulting persistence diagrams follows from the functoriality of the construction and elementary interleaving arguments.
@@
  • We employ crosscut complexes to pass from posets to simplicial complexes while preserving homotopy type.
@@
  • We prove stability of the resulting persistence diagrams under perturbations of the input metric in a density-based instantiation.
@@ 2.1 Persistent Homology
The bottleneck distance $d_{B}$ between persistence diagrams is defined as the infimum over all matchings between diagrams of the maximum $L^{\infty}$-distance between matched points.
@@ 2.2 Finite Topological Spaces and Posets
Finite topological spaces admit a convenient combinatorial description in terms of preorders and posets. We briefly recall the basic constructions and results that will be used later; a comprehensive reference is [1].
@@
While order complexes may be large, they often admit smaller simplicial models with the same homotopy type. Let $\mathcal{P}$ be a finite poset. A subset $\mathcal{C}\subseteq\mathcal{P}$ is called a crosscut if:
  1. (i)
    $\mathcal{C}$ is an antichain;
  2. (ii)
    every chain in $\mathcal{P}$ is comparable with at least one element of $\mathcal{C}$;
  3. (iii)
    for any $\mathcal{A}\subseteq\mathcal{C}$, if $\mathcal{A}$ has an upper bound or a lower bound in $\mathcal{P}$, then the join $\bigvee\mathcal{A}$ or the meet $\bigwedge\mathcal{A}$ exists in $\mathcal{P}$ (see [2] for details).
Definition.
Let $\mathcal{C}$ be a crosscut in a finite poset $\mathcal{P}$. The crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ is the simplicial complex whose simplices are the subsets of $\mathcal{C}$ that admit a common upper bound or a common lower bound in $\mathcal{P}$.
Theorem 5 (Crosscut Theorem [6, Theorem 3], [2, Theorem 10.8]).
The crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ is homotopy equivalent to the order complex $\Delta(\mathcal{P})$.
@@ 3.1 Topological Models from Metric Data
From $\mathcal{P}$ we may form its order complex $\Delta(\mathcal{P})$, as well as crosscut complexes associated to suitable crosscuts.
Proposition 6.
Let $(X,\mathcal{T})$ be a finite topological space, $\mathcal{P}$ the associated poset, and $\mathcal{C}$ a crosscut consisting of maximal elements of $\mathcal{P}$. Then $(X,\mathcal{T})$ and the crosscut complex $\Delta(\mathcal{P},\mathcal{C})$ are weakly homotopy equivalent.
Proof.
By Theorem 3, $(X,\mathcal{T})$ is homotopy equivalent to its $T_{0}$-quotient. By Theorem 4, this quotient is weakly homotopy equivalent to the order complex $\Delta(\mathcal{P})$. The result now follows from the Crosscut Theorem (Theorem 5). ∎
Thus, any topology on $X$ determines a simplicial complex that is typically much smaller and more amenable to homological computation.
@@ 3.2 Filtrations and Functoriality
$\mathbb{K}\colon I\longrightarrow\mathbf{Simp},\qquad i\longmapsto K_{i}:=\Delta(\mathcal{P}_{i},\mathcal{C}_{i}).$
@@
Each step of the construction—passing from a finite topological space to its associated poset, and from a poset to a crosscut complex—is functorial with respect to continuous maps. Moreover, the weak homotopy equivalences involved are natural with respect to these maps. Applying homology yields naturally isomorphic persistence modules. In particular, the induced maps on homology commute with the structure maps of the filtrations. ∎
@@
Before forming the crosscut complexes, one may further simplify the posets $\mathcal{P}_{i}$ using standard techniques from poset combinatorics, such as removal of beat points or other homotopy-preserving reductions. Since these simplifications do not change the homotopy type, the resulting persistence diagrams remain unchanged by the Persistence Equivalence Theorem (Theorem 2).
@@
3.3 A Density-Based Filtration
We conclude this section by describing a concrete instance of the above framework. Let $t_{1}\leq t_{2}\leq\cdots\leq t_{m}$ be real parameters. For each $i$, let $\mathcal{O}_{i}$ be a collection of subsets of $X$ determined by a density criterion with threshold $t_{i}$, and let $\mathcal{T}_{i}$ be the topology generated by $\mathcal{O}_{i}$. The monotonicity of the density thresholds ensures that
@@ 3.3 A Density-Based Filtration
In the extreme cases, $t_{1}$ may be chosen sufficiently small so that all co-singleton subsets are declared open, while $t_{m}$ may be chosen sufficiently large so that $\mathcal{T}_{m}$ is the trivial topology. Accordingly, the associated crosscut complexes interpolate between a discrete set of vertices and a single vertex.
Stability properties of this construction under perturbations of the metric data will be established in Section 4.
@@ 4 Stability
In this section we establish stability of the persistence modules arising from our construction under perturbations of the input metric. The proof relies on the functorial nature of the construction and on the theory of interleavings recalled in Section 2.1.

### v2

@@ Abstract
We propose a functorial framework for persistent homology based on finite topological spaces and their associated posets. Starting from a finite metric space, we associate a filtration of finite topologies whose structure maps are continuous identity maps. By passing functorially to posets and to order complexes, we obtain persistence modules without requiring inclusion relations between the resulting complexes. We show that standard poset-level simplifications preserve persistent invariants and establish stability of the resulting persistence diagrams under perturbations of the input metric in a basic density-based instantiation, illustrating how stability arguments arise naturally in our framework. We further introduce a concrete density-guided construction, designed to be faithful to anchor neighborhood structure at each scale, and demonstrate its practical viability through an implementation tested on real datasets.
@@
Keywords: Persistent homology; finite topological spaces; posets; order complexes; stability.
@@ 1 Introduction
In this paper, we develop a functorial framework for persistent homology that interpolates between metric data and simplicial complexes by way of finite topological spaces and posets. Starting from a finite metric space, we first associate to it a family of finite topologies indexed by a scale parameter. These topologies encode relational or density-based information and form a filtration via continuous identity maps. By passing functorially to associated posets and then to order complexes, we obtain persistence modules without requiring inclusion relations between the resulting complexes.
This perspective separates the role of the metric as a generator of topological structure from the subsequent homological analysis. As a result, it allows for topological and combinatorial simplifications at the poset level, while preserving persistent invariants. Moreover, for a basic density-based instantiation of the framework, stability of the resulting persistence diagrams follows from functoriality and elementary interleaving arguments. This toy case serves to illustrate how stability considerations arise naturally in our setting, even though it does not yet capture informative structural features of the data.
In particular, in the basic density-based filtration the associated posets become contractible as soon as uncovered points appear, causing the topology to collapse to a cone and limiting the descriptive power of the resulting persistence. This observation motivates the introduction of a more refined, density-guided construction that aims to remain faithful to anchor neighborhood structure throughout the filtration.
@@
  • We establish stability of the resulting persistence diagrams under perturbations of the input metric in a basic density-based instantiation, illustrating the natural compatibility of the framework with interleaving-based stability arguments.
  • We introduce a concrete density-guided construction designed to preserve anchor neighborhood structure across scales, and provide an implementation demonstrating its applicability to real datasets.
@@ 2.1 Persistent Homology
The bottleneck distance $d_{B}$ between persistence diagrams is defined as the infimum over all matchings between diagrams of the maximum $\ell^{\infty}$-distance between matched points.
@@ 2.2 Finite Topological Spaces and Posets
Finite topological spaces admit a convenient combinatorial description in terms of preorders and posets. We briefly recall the basic constructions and results that will be used later; see Barmak [1] for a comprehensive modern treatment, building on earlier work of Alexandrov and McCord.
@@ 3.1 Topological Models from Metric Data
From $\mathcal{P}$ we may form its order complex $\Delta(\mathcal{P})$. By Theorem 3 and Theorem 4, the finite topological space $(X,\mathcal{T})$ is weakly homotopy equivalent to the order complex $\Delta(\mathcal{P})$. Thus, any topology on $X$ determines a simplicial complex in a canonical way and standard computational tools for simplicial complexes become available for homology calculations.
@@ 3.2 Filtrations and Functoriality
$\mathbb{K}\colon I\longrightarrow\mathbf{Simp},\qquad i\longmapsto K_{i}:=\Delta(\mathcal{P}_{i}).$
@@
Each step of the construction—passing from a finite topological space to its associated poset, and from a poset to its order complex—is functorial with respect to continuous maps. Moreover, the weak homotopy equivalences involved are natural with respect to these maps. Applying homology yields naturally isomorphic persistence modules. In particular, the induced maps on homology commute with the structure maps of the filtrations. ∎
@@
Before forming the order complexes, one may further simplify the posets $\mathcal{P}_{i}$ using standard techniques from poset combinatorics, such as removal of beat points or other homotopy-preserving reductions. Since these simplifications do not change the homotopy type, the resulting persistence diagrams remain unchanged by the Persistence Equivalence Theorem (Theorem 2).
@@
3.3 Density-Based Filtration – Basic Model
We conclude this section by describing a toy case application of the above framework. Let $t_{1}\leq t_{2}\leq\cdots\leq t_{m}$ be real parameters. For each $i$, let $\mathcal{O}_{i}$ be a collection of subsets of $X$ determined by a density criterion with threshold $t_{i}$, and let $\mathcal{T}_{i}$ be the topology generated by $\mathcal{O}_{i}$. The monotonicity of the density thresholds ensures that
@@ 3.3 Density-Based Filtration – Basic Model
In the extreme cases, $t_{1}$ may be chosen sufficiently small so that all co-singleton subsets are declared open, while $t_{m}$ may be chosen sufficiently large so that $\mathcal{T}_{m}$ is the trivial topology. Accordingly, the associated order complexes interpolate between a discrete set of vertices and a single vertex.
Despite its conceptual clarity, this basic construction has a significant limitation: as soon as uncovered points (i.e. points not contained in any of the chosen generating neighborhoods at that scale) appear, the associated posets acquire a maximal element, causing the topology to become contractible. Consequently, the filtration rapidly loses higher-order information about the data. While this behavior makes the basic model unsuitable for detailed data analysis, it serves as a useful conceptual setting in which functoriality and stability can be examined transparently.
Stability properties of this basic construction under perturbations of the metric data will be established in Section 4, serving to exemplify how stability arguments arise naturally within our functorial framework.
@@ 4 Stability
In this section we establish stability of the persistence modules arising from the basic density-based construction under perturbations of the input metric. The proof relies on the functorial nature of the construction and on the theory of interleavings recalled in Section 2.1.
@@ 4.3 Stability of Persistence Diagrams
5 Density-Guided Instance
The basic density-based filtration introduced in Section 3.3 serves as a conceptual illustration of the functorial framework and its stability properties. As discussed there, however, that construction rapidly becomes topologically trivial once the topology admits a maximal element, limiting its descriptive power for real data. In this section, we introduce a concrete density-guided construction designed to overcome this limitation and report on a preliminary empirical study based on an explicit implementation.
The purpose of this section is not to provide theoretical guarantees or to interpret empirical results in depth, but rather to demonstrate that the proposed framework admits practical instantiations that produce informative persistence modules on real datasets.
5.1 The density-guided construction
Let $(X,d)$ be a finite metric space. As before, we consider a finite sequence of density thresholds
$\tau_{1}\leq\tau_{2}\leq\cdots\leq\tau_{m},$
and associate to each $\tau_{i}$ a finite topology $\mathcal{T}_{i}$ on the underlying set $X$.
In the basic density-based construction of Section 3.3, the topology $\mathcal{T}_{i}$ is generated by declaring all subsets of $X$ satisfying a density condition at scale $\tau_{i}$ to be open. While conceptually simple, this approach is computationally infeasible in practice, as the number of such dense subsets grows exponentially with $|X|$.
The density-guided construction replaces this exhaustive approach with a structured and tractable generator set. We begin by fixing a finite collection of candidate anchor neighborhoods
$\mathcal{A}=\{A_{1},\dots,A_{k}\},$
determined once from the metric data. These anchor neighborhoods serve as potential generators of the topology across all scales.
For each threshold $\tau_{i}$, only a subset
$\mathcal{A}_{i}\subseteq\mathcal{A}$
of anchor neighborhoods satisfies the prescribed density criterion at scale $\tau_{i}$. As the density threshold increases, fewer anchor neighborhoods remain eligible, yielding a nested sequence
$\mathcal{A}_{1}\supseteq\mathcal{A}_{2}\supseteq\cdots\supseteq\mathcal{A}_{m}.$
To control the evolution of the topology and preserve relational information between dense regions, we augment the eligible anchor neighborhoods $\mathcal{A}_{i}$ by a secondary collection of sets $\mathcal{B}_{i}$. The sets in $\mathcal{B}_{i}$ are not density generators themselves, but are introduced to ensure compatibility of the topology across successive scales and to prevent loss of structural information when anchor neighborhoods become ineligible. Their construction is constrained by the topology at the previous scale and by the requirement that the resulting topology remain faithful to the anchor structure. The topology $\mathcal{T}_{i}$ is then generated by the family
$\mathcal{G}_{i}:=\mathcal{A}_{i}\cup\mathcal{B}_{i}.$
Remark.
While the explicit construction of the sets $\mathcal{B}_{i}$ involves several layers and is deferred to the implementation, their role is purely structural: they serve to mediate transitions between successive anchor configurations without introducing new density-driven features.
The defining feature of the construction is that $\mathcal{T}_{i}$ is faithful to anchor structure in the following sense: every minimal nonempty intersection of anchor neighborhoods in $\mathcal{A}_{i}$ appears as an element of the minimal basis of $\mathcal{T}_{i}$. Equivalently, the collection of minimal nonempty intersections generated by $\mathcal{A}_{i}$ is contained in the minimal basis of the topology generated by $\mathcal{G}_{i}$.
As the density threshold increases, the shrinking of the eligible anchor set $\mathcal{A}_{i}$ drives the evolution of the topology across scales and prevents premature collapse to contractible structures. The resulting family of topologies
forms a filtration via continuous identity maps, to which the functorial constructions of Sections 3.1 and 3.2 apply.
Remark.
The density-guided construction may thus be viewed as a filtration obtained by progressively invalidating anchor neighborhoods rather than introducing new generators.
5.2 Implementation
The density-guided construction described above has been implemented in a research prototype software pipeline. The implementation follows the functorial structure described in Sections 3.1 and 3.2 and produces, for each scale parameter, the associated finite topology, its corresponding $T_{0}$-poset, and the resulting order complex. Homology is computed on the $2$-skeleton of the order complex, which is sufficient for dimensions 0 and $1$.
In addition to computing persistence modules, the implementation generates detailed diagnostic outputs at each filtration step, including per-scale summaries of topological and combinatorial quantities, persistence barcodes in both visual and machine-readable formats, and auxiliary plots for inspection. These outputs are included directly in the repository so that the empirical behavior of the construction can be examined without re-running the experiments.
The implementation used for the empirical results reported in this section is publicly available [4].
5.3 Datasets
We evaluated the implementation on two datasets commonly used in topological data analysis and related areas:
  • COIL-20, a collection of grayscale images of objects observed from multiple viewpoints [5].
  • Paul15, a high-dimensional dataset frequently used in empirical studies of persistent homology [7].
These datasets were chosen solely to verify that the construction produces well-defined filtrations and computable persistence modules on real data, and not for comparative evaluation.
In addition, the repository includes several synthetic examples designed to illustrate and probe the behavior of the density-guided construction under controlled conditions. These examples are intended to provide intuition about the evolution of the associated topologies and posets across density scales, and to serve as diagnostic test cases for the implementation.
5.4 Empirical observations
For both datasets, the density-guided construction produced nontrivial filtrations of finite topological spaces whose associated posets retained meaningful structure across multiple scales. In particular, the resulting order complexes did not collapse immediately to contractible spaces as the density threshold increased.
Using the implementation, we compute summary tables and persistence diagrams for homological dimensions 0 and $1$, documenting the evolution of connected components and one-dimensional features across density scales. The repository includes per-step filtration summaries (recording, among other quantities, Betti numbers, numbers of anchors, coverage statistics, and combinatorial sizes of the associated complexes), as well as barcode visualizations and machine-readable barcode data.
These results are presented descriptively and without further interpretation. No claims are made regarding semantic interpretation of features or comparative performance; the purpose of this section is solely to document the empirical behavior of the construction.
5.5 Remarks
We conclude this section with several remarks intended to place the density-guided construction in context.
First, while density plays a central role in many data analysis methods, persistent homology pipelines in which the scale parameter itself is driven by density appear to be relatively unexplored. The present framework provides a natural setting for studying how dense regions interact and merge across density levels.
Second, the construction is not intended for geometric shape recovery. Instead, it aims to capture relational and statistical structure across density scales, a task that is orthogonal to classical goals in persistent homology but of independent interest.
Finally, an advantage of the proposed approach is that homological features can be readily traced back to subsets of data points through the underlying poset structure. This makes it straightforward to identify the supports of homology classes, a task that is often nontrivial in standard matrix-reduction-based pipelines. A detailed investigation of computational aspects and theoretical properties of this feature is left for future work.